### v1

# Thermal circuit model for silicon quantum-dot array structures

Abstract
Temperature rise of qubits due to heating is a critical issue in large-scale quantum computers based on quantum-dot (QD) arrays. This leads to shorter coherence times, induced readout errors, and increased charge noise. Here, we propose a simple thermal circuit model to describe the heating effect on silicon QD array structures. Noting that the QD array is a periodic structure, we represent it as a thermal distributed-element circuit, forming a thermal transmission line. We validate this model by measuring the electron temperature in a QD array device using Coulomb blockade thermometry, finding that the model effectively reproduces experimental results. This simple and scalable model can be used to develop the thermal design of large-scale silicon-based quantum computers.
I Introduction
Silicon quantum-dot (QD) arrays are promising candidates for scalable quantum computing platforms because of their outstanding transistor integration, long coherence time, and high fidelity Burkard et al. (2023); Kodera (2024); Vandersypen et al. (2017); Veldhorst et al. (2017); Borsoi et al. (2024); Li et al. (2018); Philips et al. (2022); Lee et al. (2020, 2022). For realizing a large-scale integration, heating in the QD array is an unavoidable issue. The rise in electron (or hole) temperature due to heating leads to shorter coherence times, induced readout errors, and increased charge noise Huang et al. (2024); Kodera (2024). There are several heat sources (e.g., heat and thermal noise inflows through wires, loss of microwave signals, and driving in charge sensors) that are expected to escalate with the large-scale integration of qubits Krinner et al. (2019); Savin et al. (2006); Kawakami et al. (2013); Takeda et al. (2018); Noah et al. (2024); Undseth et al. (2023).
Two approaches have been investigated to address the heating problem. One is to develop cryo-electronics, effectively reducing the wiring from room temperature and reducing the heat input from the wiring Xue et al. (2021); Pauka et al. (2021); Bartee et al. (2024). The other is operating qubits at high temperature using heat-tolerant techniques, such as a Pauli spin blockade readout, which has been demonstrated to operate at temperatures as high as one kelvin Petit et al. (2018); Ono, Mori, and Moriyama (2019); Petit et al. (2020); Yang et al. (2020); Petit et al. (2022); Camenzind et al. (2022); Huang et al. (2024). Meanwhile, we are exploring another approach, namely thermal management through the design of heat inflow paths in QD array structures and the control of heat sources using real-time ambient temperature measurements. To achieve this, the thermal conduction characteristics of the device need to be understood. To accurately measure the thermal characteristics, local heaters and thermometers must be integrated into the QD array.
Coulomb blockade thermometry (CBT) is known to measure temperature in QD structures in cryogenic environments Beenakker (1991); Giazotto et al. (2006); Rossi, Ferrus, and Williams (2012). CBT has been used to study the thermal conduction properties of quantum point contacts (QPCs) Molenkamp et al. (1992); Chiatti et al. (2006) and single-electron transistors (SETs) Hoffmann et al. (2007, 2009); Dutta et al. (2017) in cryogenic environments. However, the thermal conduction characteristics of QD array structures have yet to be investigated. Recently, thermal analysis towards silicon quantum computers has been conducted. Heating in cryo-electronic circuits was investigated using two transistors for the local heater and CBT sensor de Kruijf et al. (2024), and thermal transient in QD array structure was obtained by using microwave pulses and reflectometry Champain et al. (2024). Furthermore, the thermal circuit model based on the lumped-element model was developed to analyze the cryogenic CMOS on-chip thermometry, a crucial step for understanding the thermal characteristics in cryo-chips Noah et al. (2024).
In this paper, to clarify the thermal conduction characteristics of QD array structures, we propose a simple thermal circuit model. In Sec. II, we explain our model, where the QD array structure is described by an effective thermal circuit based on the distributed-element model. This results in a analytical expression of the thermal transmission line. In Sec. III, we show the experimental validation of our model, where we measure the electron temperature in a QD array device using CBT. A local heater is implemented by flowing current through the barrier gates of the QD array. This experimental setup mimics the heat generated by the local current flowing through the gate wiring for the qubit addressing Kawakami et al. (2013); Li et al. (2018) and by the microwaves applied to the gate wiring for an electric dipole spin resonance Burkard et al. (2023). From the distance dependence between the local heater and the SET, we find that this model reproduces the experimental results. In Sec. IV, we discuss the limitations of our model and future work. Our proposed model is simple, intuitive, and scalable and can be used for the thermal management of large-scale quantum computers in silicon.
II Model
Figure 1: Conceptual diagram of the heating effects in silicon quantum computers. Heat flow is modeled by describing the electron temperature $T_{e}$ of qubits as a function of the heating power $P$ and various parameters that depend on the device structure.
We describe the heating effects in silicon quantum computers separately in three parts: (a) Heating source, (b) heat flow, and (c) qubits. Figure 1 illustrates an conceptual diagram. In silicon quantum computers that are typically implemented in cryogenic environments using dilution refrigerators, there are several heat sources such as radiation within dilution refrigerator, heat and thermal noise inflows through wires, loss of microwave signals, and local currents in charge sensors Krinner et al. (2019); Savin et al. (2006); Kawakami et al. (2013); Takeda et al. (2018); Noah et al. (2024); Undseth et al. (2023). These heat sources can be modeled as heating power $P$. The heat inflow to the qubits and the heat dissipation to the cooler depend on the device structure, specifically its thermal conductance and capacitance. Our objective is to model the electron temperature of qubits $T_{e}$ as a function of $P$ and various parameters depending on the device structure. By employing this model, we can estimate $T_{e}$ from $P$ to design the heat flow path accordingly.
For constructing the thermal model of the silicon QD array, we note that typical silicon QD array structures are configured with a periodic gate array Vandersypen et al. (2017); Veldhorst et al. (2017); Borsoi et al. (2024); Li et al. (2018); Philips et al. (2022); Lee et al. (2020, 2022), and this periodicity is essential for modeling their thermal characteristics. In our model, the QD array device is represented as a periodic structure shown in Fig. 2(a). The unit cell comprises a set of a plunger gate (PG), barrier gate (BG), and insulators such as SiO2, where the length of the unit cell in $x$ direction is $L_{\rm{cell}}$. For simplicity, the silicon channel and the regions outside the gate structure are omitted. Heat flow is considered two-dimensional due to the presence of thick SiO2 or Si layers at the top and bottom, which have lower thermal conductivity than the gate structures made of metal or polysilicon at cryogenic temperatures. This heat flow is divided into two paths: heat inflow path and dissipation path. The heat inflow path represents the propagation of heat from the source to the qubits along the periodic gate structure, while the dissipation path represents the heat propagation from the heat source to the cooler. Here, the heat flow that does not reach the qubits is dissipated through a medium with high thermal conductivity, such as metallic gates, metallic wires, and metallic structure in a dilution refrigerator.
Figure 2: (a) Model of the silicon QD array structure. The periodic gate structure comprises the plunger gate (PG) and the barrier gate (BG) made of metal or polysilicon. Top and bottom layers are thick Si or SiO2. Heat inflow propagates to the $x$ direction, and heat dissipation propagates to the other directions. (b) Simplified thermal rod model for periodic structure in (a) and the thermal circuit analogous to the transmission line in an electric circuit.
In our model, the thermal structure of the QD array is simplified as a thermal rod, as shown in Fig. 2(b). We employ a thermal circuit model where thermal characteristics are analogously represented to an electrical circuit. In this model, heating power and temperature are represented by current and voltage, respectively, similar to an electrical circuit model Noah et al. (2024). Thus, periodic structures can be modeled using a thermal distributed-element model, analogous to the transmission line model in electrical circuits. We show the thermal circuit diagram for the thermal rod in Fig. 2(b), where unlike in the case of electronic circuits, inductors are excluded. This thermal circuit can be analyzed using the telegrapher’s equations. In this context, $R_{\rm{in}}$ (heat inflow resistance), $R_{\rm{ex}}$ (heat dissipation resistance), and $C_{\rm{ex}}$ (heat capacitance) are defined as distributed elements, as shown in Fig. 2(b).
Here, we outline the assumptions underlying our model. (i) The first assumption is that the thermal circuit consists of an infinite periodic structure. This assumption is valid for large-scale QD array structures and thermally uniform structures such as those incorporating dummy metals in the gate layer. (ii) The second assumption is that the distributed elements of the thermal circuit, $R_{\rm{in}}$, $R_{\rm{ex}}$, and $C_{\rm{ex}}$, are temperature dependent and spatially uniform. To account for the temperature dependence of these spatially uniform circuit elements, we consider the device’s net effective temperature, which is the spatially averaged temperature in the area to be analyzed. (iii) The third assumption is that this net effective temperature is proportional to $T_{e}$. This is an assumption that is also applied in Ref. Noah et al. (2024). These assumptions allow us to use a transmission line model based on the distributed-element model. While the distributed circuit elements are spatially uniform, the spatial temperature distribution of the QD array is not uniform and decays exponentially [as derived in following Eq. (4)]. The approximation error arising from this inconsistency is minimized when the spatial temperature variation is smaller than the net effective temperature.
We then explain the thermal circuit elements $R_{\rm{in}}$, $R_{\rm{ex}}$, and $C_{\rm{ex}}$, respectively. $R_{\rm{in}}$ is the thermal resistance (per unit length) for heat inflow from the heat source to the temperature measurement point (qubit). This thermal resistance is a series of resistances of a periodic structure, which is composed of the metallic or polysilicon gates and the insulator (we show the order estimation of $R_{\rm{in}}$, see Appendix C). On the basis of the above discussion, we use $R_{\rm{in}}\propto 1/T_{e}^{\beta_{\rm{in}}}$, where $\beta_{\rm{in}}$ depends on the thermal conductivity in cryogenic regime and the thickness of the gate and insulator. For example, $\beta_{\rm{in}}\approx 1$ when the gate is dominant and $\beta_{\rm{in}}\approx 2$ when the insulator (amorphous SiO2) is dominant (see Appendix A).
$R_{\rm{ex}}$ is the thermal resistance (inverse of the thermal conductance per unit length) for heat dissipation from the heat source. Heat is dissipated through metallic wires and metallic structures in a dilution refrigerator and is finally collected in the mixing chamber, which acts as a cooler. Therefore, $R_{\rm{ex}}$ depends on the structure of the whole chip and chip’s implementation form such as the circuit board mounting the chip, bonding wires, sample packaging, and cables. Most of these materials are metallic. We then use $R_{\rm{ex}}\propto 1/T_{e}^{\beta_{\rm{ex}}}$, where $\beta_{\rm{ex}}\approx 1$ considering the metallic materials are dominant (see Appendix A).
The heat capacitance $C_{\rm{ex}}$ is crucial for real-time thermal management because it influences the temporal changes in device temperature, which is essential for understanding the dynamic temperature effects on qubits, such as heat-induced frequency shifts Takeda et al. (2018); Undseth et al. (2023); Sato and Kawahara (2024). It is noted that specific heat changes drastically (approximately $\propto T^{3}$) at cryogenic temperatures, according to the Debye model, as shown in Appendix A.
In analogy to transmission line models in electrical circuits, $T_{e}$, the temperature of the thermal rod at a point that is a distance $D$ away from the heat source having heating power $P$, can be derived as follows. From the telegrapher’s equations, we obtain the relations ${T_{e}(x+D)=T_{e}(x)e^{-\gamma_{\rm{th}}D}}$, ${P(x+D)=P(x)e^{-\gamma_{\rm{th}}D}}$, and ${T_{e}(x)=Z_{\rm{th}}P(x)}$, where the position on the thermal rod $x$, thermal characteristic impedance $Z_{\rm{th}}$ and thermal propagation constant $\gamma_{\rm{th}}$ are introduced. From these relations, we can describe
$T_{e}(x+D)=Z_{\rm{th}}e^{-\gamma_{\rm{th}}D}P(x).$ (1)
Since $x$ is arbitrary, we will denote $T_{e}(x+D)$ simply as $T_{e}$ and $P(x)$ as $P$ from here on.
We then describe $Z_{\rm{th}}$ and $\gamma_{\rm{th}}$ as
$Z_{\rm{th}}$ $=\sqrt{\frac{R_{\rm{in}}}{\left(\frac{1}{R_{\rm{ex}}}+j\omega C_{\rm{ex}}\right)}}\approx\sqrt{R_{\rm{in}}R_{\rm{ex}}}\approx a^{1+\beta}/T_{e}^{\beta},$ (2)
$\gamma_{\rm{th}}$ $=\sqrt{R_{\rm{in}}\left(\frac{1}{R_{\rm{ex}}}+j\omega C_{\rm{ex}}\right)}\approx\sqrt{\frac{R_{\rm{in}}}{R_{\rm{ex}}}},$ (3)
where $j$ and $\omega$ are the imaginary unit and angular frequency of thermal behaviors, respectively. The terms $a$ and $\beta$ are constant parameters depending on the device, where $\beta=(\beta_{\rm{in}}+\beta_{\rm{ex}})/2$. In the first approximation in Eqs. (2) and (3), $\omega$ is set to zero to perform the static analysis. The second approximation in Eq. (2) is derived by the relations $R_{\rm{in}}\propto 1/T_{e}^{\beta_{\rm{in}}}$ and $R_{\rm{ex}}\propto 1/T_{e}^{\beta_{\rm{ex}}}$.
Substituting Eq. (2) into Eq. (1), we obtain
$T_{e}\approx ae^{-\frac{\gamma_{\rm{th}}}{1+\beta}D}P^{\frac{1}{1+\beta}}.$ (4)
This result indicates that the electron temperature is modeled as a function of heating power $P$, the distance $D$, and three device-dependent parameters ($a$, $\gamma_{\rm{th}}$, and $\beta$).
Hereafter, for simplicity, we use $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\approx 1$, i.e., $\beta\approx 1$. This approximation is valid because, in a typical QD array structure, the thickness of the insulator is relatively lower than that of the gate, justifying $\beta_{\rm{in}}\approx 1$. Similarly, $\beta_{\rm{ex}}\approx 1$ is valid considering that metallic materials dominate the heat dissipation path, as mentioned above. Therefore, we can rewrite Eq. (4) as
$T_{e}\approx ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P}.$ (5)
We note that $a$ depends on the multiple of the thermal resistances ($R_{\rm{in}}\times R_{\rm{ex}}$) from Eq. (2). On the other hand, $\gamma_{\rm{th}}$ depends on the ratio of the thermal resistances ($R_{\rm{in}}/R_{\rm{ex}}$) from Eq. (3). Furthermore, both $a$ and $\gamma_{\rm{th}}$ are independent of $T_{e}$. The term $2/\gamma_{\rm{th}}$ represents the characteristic length of the heat dissipation, indicating that the heating effect decreases by $1/e$. Thereby, the temperature independent parameters $a$ and $\gamma_{\rm{th}}$ thermally characterize the QD array structure.
Here, we extend our model to the case of multiple heat sources. Typically, the electron temperature is written as $T_{e}^{\prime}=\sqrt{T_{\rm{MXC}}^{2}+T_{e}^{2}}$, where $T_{e}^{\prime}$, $T_{\rm{MXC}}$, and $T_{e}$ are the measurable temperature, the base temperature of refrigerator, and electron temperature, respectively Noah et al. (2024); de Kruijf et al. (2024). Consistent with this notation, we employ the following model, i.e., the total rise in electron temperature $T_{e}^{\rm{total}}$ due to multiple heat sources is calculated using the root sum square as
$T_{e}^{\rm{total}}=\sqrt{\sum_{i=1}^{N}(\Delta T_{e}^{i})^{2}},$ (6)
where $N$ is the number of heat sources and $\Delta T_{e}^{i}$ is the rise in electron temperature due to the $i$-th heat source. From Eq. (6), Eq. (5) can be rewritten under the background temperature $T_{0}$ as
$T_{e}^{\rm{total}}=\sqrt{T_{e}^{2}+T_{0}^{2}}=ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P+P_{0}},$ (7)
where $T_{0}=ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P_{0}}$, and $P_{0}$ is the background heating power. We use this equation to fit experimental results. Note that in our experiment, $T_{0}$ and $P_{0}$ includes the effect of lifetime broadening and charge noise broadening de Kruijf et al. (2024).
III Experiment
III.1 Device structure
To verify the proposed thermal model, we experimentally measure $T_{e}$ in a QD array. Figure 3(a) shows the device structure Utsugi et al. (2023). In this device, we can measure the temperature at different distances from the heating point. The T-shaped silicon channel (green) is fabricated using silicon on insulator (SOI), and multiple polysilicon gate electrodes are formed on the T-shaped Si channel. The horizontal channel is covered with first gates (BG0-BG3, light blue) and second gates (PG0-PG3, blue), for a total of seven gates. These fine-pitch gate structures are fabricated using the self-align patterning process Lee et al. (2020); kun . Each gate has terminals on both sides, allowing current to flow through them by applying DC voltages through wiring connected to instruments at room temperature.
Another second gate (SGS, red) and other third gates (TG1 and TG2, orange) on the vertical channel comprise a single-electron transistor (SET), and temperature around the SET can be measured using CBT as described below. In this experiment, we treat the temperature measured by the SET as the electron temperature $T_{e}$.
The four first gates (BG0-BG3) are used as heat sources by flowing current through each gate. Figure 3(b) shows the cross-section between A and A’ in Fig. 3(a). Corresponding to the model shown in Fig. 2(a), the unit cell length of this device is $L_{\rm{cell}}=120$ nm (we also measure the device of $L_{\rm{cell}}=160$ nm, see Appendix B).
The silicon chip with the above structure is glued to a cryogenic printed circuit board (QBoard, Qdevil) using silver paste, and each terminal is wired using aluminum bonding wire. Measurements were performed using a dilution refrigerator (Proteox, Oxford Instruments) with a base temperature of 8 mK.
III.2 Measurement
We use CBT to estimate the local temperature at the SET. In CBT, there are several restrictions for QD parameters for accurately measuring temperature, e.g., the tunnel rate of barriers of SET $h\Gamma$, energy spacing of the quantum levels $\Delta E$, thermal energy $k_{B}T$, and charging energy $e^{2}/C$  Beenakker (1991). In this experiment, we use the classical regime, i.e., $h\Gamma\ll\Delta E\ll k_{B}T\ll e^{2}/C$. By measuring the current $I_{d}$ flowing through the SET while sweeping the voltage of the SGS gate $V_{g}$ and the drain $V_{d}$ [Fig. 4(a)], we can obtain the SGS gate voltage dependence of the differential conductance $G=dI_{d}/dV_{d}$ [Fig. 4(b)] and estimate the electron temperature around SET $T_{e}$. The differential conductance $G$ in the region where $V_{d}$ is small is Beenakker (1991)
$\frac{G}{G_{\rm{max}}}=\frac{\alpha e(V_{g}-V_{0})}{k_{B}T_{e}\sinh\left(\frac{\alpha e(V_{g}-V_{0})}{k_{B}T_{e}}\right)}\approx\cosh^{-2}\left(\frac{\alpha e(V_{g}-V_{0})}{2.5k_{B}T_{e}}\right),$ (8)
where $G_{\rm{max}}$, $\alpha$, $e$, $k_{B}$, and $V_{0}$ are the peak value of $G$, the lever arm, the elementary charge, the Boltzmann constant, and the voltage value at $G_{\rm{max}}$, respectively. The lever arm was estimated as $\alpha=0.08$ by obtaining the Coulomb diamond shown in Fig. 4(a). Figure 4(c) and 4(d) show examples of $G$ (Coulomb peak) at different $T_{e}$ situations. It can be confirmed that the Coulomb peak shape reflects the change in $T_{e}$, and $T_{e}$ can be estimated by fitting the experimental result by the curve of Eq. (8). This measurement method is constrained by the tunnel rate and can be used in the range $k_{B}T_{e}\gg h\Gamma$, where $h\Gamma/k_{B}\sim 1$ K in this experiment.
Figure 3: (a) Schematic diagram of the device structure and setup for electron temperature measurement. An SOI channel (green), BGs (light blue), TGs (orange), and SGS (red) have electrodes to apply voltages, where each wiring is a twisted cable made of phosphor bronze and has a low pass filter mounted on Qboard. (b) Schematic diagram of the cross-section at A-A’ in (a). The BGs made of polysilicon are used as a heater by flowing current. Gate SGS is used as the plunger gate in SET.
We use a local heater by applying a current to each of the four gate electrodes BG0-BG3 as shown in Fig. 5(a). If the electric field around the SET changes during the measurement, the SET conditions are affected by the electric field and the temperature measurement cannot be performed correctly. Therefore, we apply a voltage of $+V/2$ and $-V/2$, where $V=\sqrt{RP}$ ($R$ is the resistance of the entire gate wiring) to each terminal of the gate to allow current to flow through the gate. Due to the symmetry of the wiring structure in the dilution refrigerator, we set the voltage to be approximately zero on the channel. This stabilizes the SET condition because the voltage condition around the SET is stable even if the heating power $P$ is changed by varying $V$, where $P=V^{2}/R$. We measure the resistance of each gate wiring excluding the additional resistance on the circuit board for a low pass filter and use these values to calculate $P$, considering that the polysilicon gate has the dominant resistance among all of the gate wiring.
Figure 4: (a,b) Coulomb diamond of (a) the drain current of SET $I_{d}$ and (b) the differential conductance of SET $G$. The (a) white and (b) black dotted lines show one of the Coulomb diamond areas. (c) and (d) show examples of Coulomb peak and fitting curve with Eq. (8) at different temperatures. The estimated temperatures are (c) 2.4 K and (d) 5.5 K.
Figure 5(b) shows the experimental results of $T_{e}$ as a function of the heating power $P$ when each gate is used as a heat source. As expected, $T_{e}$ increases monotonically with $P$ and $1/D$. The experimental results can be fitted using Eq. (7) as
$T_{e}=A_{i}\sqrt{P+P_{0}},$ (9)
[…]
We measure another device of different gate pitch as shown in Appendix B. From these results, we conclude that the proposed model successfully reproduces the experimental results. This demonstrates its capability to model the thermal conduction characteristics of silicon QD array structures effectively.
Figure 5: (a) Schematic of local heaters by flowing current in barrier gates. The resistance of each gate excluding the resistance of the low pass filter in Qboard at cryogenic temperatures are $R_{\rm{BG0}}=78.8~{}\rm{k}\Omega$, $R_{\rm{BG1}}=78.1~{}\rm{k}\Omega$, $R_{\rm{BG2}}=70.9~{}\rm{k}\Omega$ and $R_{\rm{BG3}}=61.3~{}\rm{k}\Omega$, respectively. (b) Experimental results of $T_{e}$ as a function of $P$ and $D$. The dotted curves are the best-fit results with Eq. (9). The inset shows the best-fit results of $A$ as a function of $D$ and the dotted curve is the best-fit result with $A=ae^{-\frac{\gamma_{\rm{th}}}{2}D}$, where $a=8.5~{}\rm{K}/\sqrt{\rm{\mu W}}$ and $2/\gamma_{\rm{th}}=280~{}\rm{nm}$.
IV Discussion
To consider the limitations of our model, we re-summarize the assumptions for deriving Eq. (4) according to the above discussions: (i) The thermal circuit of the model is an infinite periodic structure, (ii) $R_{\rm{in}}$ and $R_{\rm{ex}}$ are temperature dependent and spatially uniform, and (iii) the net effective temperature of the QD array is proportional to $T_{e}$. Moreover, for deriving Eq. (5), we also assume that the temperature dependence of $R_{\rm{in}}$ and $R_{\rm{ex}}$ are metallic ($\beta_{\rm{in}}\approx\beta_{\rm{ex}}\approx 1$). Despite these potential sources of modeling error, the consistency between the experimental results and the model supports the validity of our simplifications. The model validity is further investigated using the order estimation in Appendix C.
We also discuss how to design the QD array structure. To mitigate the thermal effects on the qubits, based on Eq. (5), the following conditions should be met: (I) $P$ should be reduced, (II) $D$ should be increased, and (III) the device-dependent parameters $a$ and $2/\gamma_{\rm{th}}$ should be minimized. Conditions (I) and (II) are straightforward, i.e., the amount of heating power needs to be reduced, for example, by lowering the electrical resistance of the gate, and the qubits need to be kept appropriately distant from the heat sources. Our main result is condition (III), which is useful for quantitative thermal design of the QD array structures. Reducing $a$ implies that even if the heating power is high, the temperature rise is low; in other words, the structure does not easily accumulate heat. To reduce $a\propto\sqrt[4]{R_{\rm{in}}R_{\rm{ex}}}$, the thermal resistance of the entire structure should be minimized. On the other hand, reducing $2/\gamma_{\rm{th}}$ (the characteristic length) means creating a structure that makes it difficult to transport heat to the qubits. To reduce $2/\gamma_{\rm{th}}\approx 2\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$, the ratio of heat dissipation to heat inflow should be increased. For example, an additional metallic heat sink at the top and bottom of the gate layer can be effective. This can be achieved using technologies such as through-silicon vias (TSVs) Taguchi et al. (2024).
The next challenge is to find a way to simultaneously achieve both high QD (or qubit) performance and thermal tolerance. To further comprehensively optimize the QD array structure, the effect of $T_{e}$ on the fidelity of quantum computing needs to be quantitatively estimated. The rise in $T_{e}$ is known to lead to shorter coherence times, degraded readout fidelity, and increased charge noise (low operation fidelity) Huang et al. (2024); Kodera (2024). However, these quantitative evaluations have yet to be clarified, except for certain readout fidelity Keith et al. (2019), suggesting that further experimental and theoretical studies are necessary.
In future work, we will investigate the thermal dynamic model of the QD array structure to control the heat sources using real-time ambient temperature measurements. The thermal dynamic model is also strongly related to the operation fidelity of qubits, including the heat-induced frequency shift Takeda et al. (2018); Undseth et al. (2023); Sato and Kawahara (2024). This will be modeled by the heat capacitance $C_{\rm{ex}}$ and measured by a high-speed readout method such as reflectometry Champain et al. (2024).
In conclusion, we proposed a simple thermal circuit model for the silicon QD array structure and validated our model in experiments. Our proposed model is intuitive, simple, and scalable and is applicable to the wide spread of the QD array structures for thermal analysis.
Acknowledgements.
We acknowledge N. Mertig and A. Ramsay at Hitachi Cambridge Laboratory for helpful discussions. This work is supported by JST Mooshot R&D Grant No. JPMJMS2065.
Appendix A Thermophysical properties
We summarize thermophysical properties of represented materials in the typical silicon fabrication process based on Refs. Duthil (2014); Ekin (2006); ais , as shown in Fig. 6.
Figure 6: Thermophysical properties of represented materials where $\kappa$ of metallic materials (W, Cu, Al, and Poly-Si) are extrapolated by linear function ($\propto T$) in the cryo-temperature regime Duthil (2014); Ekin (2006). Note that Si is single crystal ais , SiO2 is amorphous Ekin (2006), and Poly-Si is P-type polysilicon (B-doped = $3\times 10^{20}$ atom/cm3ais .
Appendix B Measurement result of device B
We measure another device with $L_{\rm{cell}}=160$ nm (device B), similar to the device described in the main text (device A). Figure 7 shows the results, which exhibit the same trend as those in Fig. 5. However, the electron temperature is lower than that of device A. This is attributed to the larger gate pitch, which implies a longer gate length and a higher ratio of the polysilicon gate to the insulator. This configuration likely decreases the resistance $R_{\rm{in}}$ while maintaining $R_{\rm{ex}}$, resulting in an increased thermal characteristic length $2/\gamma_{\rm{th}}\approx 2\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$. On the other hand, $a\propto\sqrt[4]{R_{\rm{in}}R_{\rm{ex}}}$ changes minimally due to the fourth root. This measurement further validates the proposed thermal model. In Appendix C, we further summarize the experimental results.
Figure 7: Experimental results of $T_{e}$ as a function of $P$ and $D$ in device B of $L_{\rm{cell}}=160$ nm. The dotted curves are the best-fit results with Eq. (9). The inset shows the best-fit results of $A$ as a function of $D$ and the dotted curve is the best-fit result with $A=ae^{-\frac{\gamma_{\rm{th}}}{2}D}$, where $a=7.0~{}\rm{K}/\sqrt{\rm{\mu W}}$ and $2/\gamma_{\rm{th}}=700~{}\rm{nm}$.
Appendix C Summary of experimental results and order estimation of heat inflow resistance
Device A Device B
$L_{\rm{cell}}$ (nm) 120 160
$a$ (K$/\sqrt{\rm{\mu W}}$) 8.5 7.0
$2/\gamma_{\rm{th}}$ (nm) 280 700
$R_{\rm{in}}T_{e}$ ($\rm{K^{2}/(m\cdot W)}$) $5.1\times 10^{14}$ $1.4\times 10^{14}$
$R_{\rm{ex}}T_{e}$ ($\rm{K^{2}\cdot m/W}$) $10$ $17$
$R^{\rm{est}}_{\rm{in}}$ ($\rm{K^{2}/(m\cdot W)}$) $4.0\times 10^{14}$ $3.5\times 10^{14}$
Table 1: The device-dependent parameters calculated from the experimental results, and order estimation results (the last row).
Table 1 presents the device-dependent parameters estimated from the experiments, where we compare devices A (main text) and B (Appendix B). To evaluate the thermal resistances, we use the relations of $R_{\rm{in}}T_{e}=a^{2}\gamma_{\rm{th}}$ and $R_{\rm{ex}}T_{e}=a^{2}/\gamma_{\rm{th}}$ derived from Eqs. (2) and (3), where $\beta=1$ is used. We analyze these results to the order estimation values of heat inflow resistance $R^{\rm{est}}_{\rm{in}}$ on the basis of the device structures and the thermophysical properties shown in Fig. 6 (see Appendix B). Note that the order estimation of $R_{\rm{ex}}$ is difficult due to the complicated heat dissipation path.
Using the model of the QD array structure shown in Fig. 2 analogous to the series resistance, we estimate $R^{\rm{est}}_{\rm{in}}$ as
$R^{\rm{est}}_{\rm{in}}=\frac{1}{SL_{\rm{cell}}}\left(\frac{L_{\rm{PolySi}}}{\kappa_{\rm{PolySi}}}+\frac{L_{\rm{SiO_{2}}}}{\kappa_{\rm{SiO_{2}}}}\right),$ (10)
where $S$ is the cross-sectional area of the heat inflow path, $L_{\rm{PolySi}}$ ($L_{\rm{SiO_{2}}}$) is the total length of the polysilicon gates BG and PG (SiO2), where $L_{\rm{cell}}=L_{\rm{PolySi}}+L_{\rm{SiO_{2}}}$, and $\kappa_{\rm{PolySi}}$ ($\kappa_{\rm{SiO_{2}}}$) is the thermal conductivity of the polysilicon gates (SiO2). We use $S=10^{-13}$ $\rm{m}^{2}$, where $S$ is calculated as the product of the width ($W$) and height ($H$) of the heat inflow path and we use $W=1~{}\rm{\mu m}$ and $H=0.1~{}\rm{\mu m}$. From the device structure, we use $L_{\rm{PolySi}}=90$ nm for device A and $L_{\rm{PolySi}}=130$ nm for device B, while $L_{\rm{SiO_{2}}}=30$ nm for both devices. From the thermophysical properties shown in Fig. 6, we use $\kappa_{\rm{PolySi}}=0.05~{}\rm{W/(m\cdot K)}$ and $\kappa_{\rm{SiO_{2}}}=0.01~{}\rm{W/(m\cdot K)}$ assuming around 1 K. We obtain the estimation results as $R^{\rm{est}}_{\rm{in}}=4.0\times 10^{14}$ for device A and $R^{\rm{est}}_{\rm{in}}=3.5\times 10^{14}$ for device B at $T_{e}=1$ K, as shown in Table 1. The lower $R_{\rm{in}}$ for device B than for device A is consistent with the experimental trend. This result demonstrates the validity of our model. Additionally, the order of $R^{\rm{est}}_{\rm{in}}$ values are comparable to the experimental results at $T_{e}=1$ K, validating the formulation in Eq. 10 at least for the order estimation. We also expect that by improving this simple model or combining it with different approaches, e.g., finite element simulations, we can develop a framework for predicting thermal characteristics more reliably and accurately.

### v2

@@
 II, we explain our model, where the QD array structure is described by an effective thermal circuit based on the distributed-element model.
-This results in a analytical expression of the thermal transmission line.
+This results in an analytical expression of the thermal transmission line.
 In Sec.
@@
 Figure 1: Conceptual diagram of the heating effects in silicon quantum computers.
-Heat flow is modeled by describing the electron temperature $T_{e}$ of qubits as a function of the heating power $P$ and various parameters that depend on the device structure.
-We describe the heating effects in silicon quantum computers separately in three parts: (a) Heating source, (b) heat flow, and (c) qubits.
+Heat flow is modeled by describing the electron temperature $T_{e}$ of qubits as a function of the heating power $P_{0}$ and various parameters that depend on the device structure.
+We describe the heating effects in silicon quantum computers separately in three parts: (a) Heat source, (b) heat flow, and (c) qubits.
 Figure 1 illustrates an conceptual diagram.
@@
 (2023).
-These heat sources can be modeled as heating power $P$.
+These heat sources can be modeled as heating power $P_{0}$.
 The heat inflow to the qubits and the heat dissipation to the cooler depend on the device structure, specifically its thermal conductance and capacitance.
-Our objective is to model the electron temperature of qubits $T_{e}$ as a function of $P$ and various parameters depending on the device structure.
-By employing this model, we can estimate $T_{e}$ from $P$ to design the heat flow path accordingly.
+Our objective is to model the electron temperature of qubits $T_{e}$ as a function of $P_{0}$ and various parameters depending on the device structure.
+By employing this model, we can estimate $T_{e}$ from $P_{0}$ to design the heat flow path accordingly.
 For constructing the thermal model of the silicon QD array, we note that typical silicon QD array structures are configured with a periodic gate array Vandersypen et al.
@@
 For simplicity, the silicon channel and the regions outside the gate structure are omitted.
-Heat flow is considered two-dimensional due to the presence of thick SiO2 or Si layers at the top and bottom, which have lower thermal conductivity than the gate structures made of metal or polysilicon at cryogenic temperatures.
-This heat flow is divided into two paths: heat inflow path and dissipation path.
+Heat is generated at a gate, and a portion of it propagates toward a qubit located some distance away.
+All of the heat is eventually dissipated outside the chip.
+Accordingly, this heat flow is divided into two paths: heat inflow path and dissipation path.
 The heat inflow path represents the propagation of heat from the source to the qubits along the periodic gate structure, while the dissipation path represents the heat propagation from the heat source to the cooler.
-Here, the heat flow that does not reach the qubits is dissipated through a medium with high thermal conductivity, such as metallic gates, metallic wires, and metallic structure in a dilution refrigerator.
+Here, the heat is dissipated mainly through a medium with relatively high thermal conductivity, such as metallic or polysilicon gates, thick SiO2 or Si at the bottom layer, metallic wires, and metallic structure in a dilution refrigerator.
 Figure 2: (a) Model of the silicon QD array structure.
@@
 Heat inflow propagates to the $x$ direction, and heat dissipation propagates to the other directions.
-(b) Simplified thermal rod model for periodic structure in (a) and the thermal circuit analogous to the transmission line in an electric circuit.
-In our model, the thermal structure of the QD array is simplified as a thermal rod, as shown in Fig.
-2(b).
+This conceptual diagram shows a case where the heat source gate and the measurement point are located at a certain distance from each other.
+(b) Thermal circuit model from the MXC to the chip in the dilution refrigerator.
+(c) Thermal circuit model for the periodic silicon QD array structure.
 We employ a thermal circuit model where thermal characteristics are analogously represented to an electrical circuit.
@@
 (2024).
-Thus, periodic structures can be modeled using a thermal distributed-element model, analogous to the transmission line model in electrical circuits.
-We show the thermal circuit diagram for the thermal rod in Fig.
-2(b), where unlike in the case of electronic circuits, inductors are excluded.
-This thermal circuit can be analyzed using the telegrapher’s equations.
-In this context, $R_{\rm{in}}$ (heat inflow resistance), $R_{\rm{ex}}$ (heat dissipation resistance), and $C_{\rm{ex}}$ (heat capacitance) are defined as distributed elements, as shown in Fig.
-2(b).
-Here, we outline the assumptions underlying our model.
-(i) The first assumption is that the thermal circuit consists of an infinite periodic structure.
-This assumption is valid for large-scale QD array structures and thermally uniform structures such as those incorporating dummy metals in the gate layer.
-(ii) The second assumption is that the distributed elements of the thermal circuit, $R_{\rm{in}}$, $R_{\rm{ex}}$, and $C_{\rm{ex}}$, are temperature dependent and spatially uniform.
-To account for the temperature dependence of these spatially uniform circuit elements, we consider the device’s net effective temperature, which is the spatially averaged temperature in the area to be analyzed.
-(iii) The third assumption is that this net effective temperature is proportional to $T_{e}$.
-This is an assumption that is also applied in Ref.
+The overall dilution refrigerator is modeled to set the base temperature of the QD array device, denoted by $T_{\rm{base}}$, as shown in Fig.
+2(b), based on Ref.
 Noah et al.
 (2024).
-These assumptions allow us to use a transmission line model based on the distributed-element model.
-While the distributed circuit elements are spatially uniform, the spatial temperature distribution of the QD array is not uniform and decays exponentially [as derived in following Eq.
-(4)].
-The approximation error arising from this inconsistency is minimized when the spatial temperature variation is smaller than the net effective temperature.
-We then explain the thermal circuit elements $R_{\rm{in}}$, $R_{\rm{ex}}$, and $C_{\rm{ex}}$, respectively.
+Assume that the mixing chamber (MXC) is cooled to a stable temperature $T_{m}$, where the cooling power and heating power are balanced.
+The effective temperature-dependent thermal resistance, denoted by $R_{m}$, includes the thermal resistance of the MXC plate, the cold finger, the printed circuit board (PCB), and wiring from the MXC plate to the chip.
+The periodic structures of the QD array can be modeled using a thermal distributed-element model, analogous to the transmission line model in electrical circuits.
+We show the thermal circuit diagram for the QD array structure in Fig.
+2(c), where unlike in the case of electronic circuits, capacitors and inductors are excluded.
+This thermal circuit can be analyzed using simultaneous differential equations analogous to the telegrapher’s equations.
+In this context, $R_{\rm{in}}$ (heat inflow resistance) and $R_{\rm{ex}}$ (heat dissipation resistance) are defined as distributed elements, as shown in Fig.
+2(c).
+The low side reference of this thermal transmission line is $T_{\rm{base}}$ and the electron temperature rise from $T_{\rm{base}}$ at position $x$ is denoted by $T_{e}(x)$.
+Here, we outline the assumptions underlying our model.
+(i) The first assumption is that the thermal circuit consists of a semi-infinite periodic structure.
+This assumption is valid for large-scale QD array structures and thermally uniform structures such as those incorporating dummy metals in the gate layer.
+(ii) The second assumption is that the distributed elements of the thermal circuit, $R_{\rm{in}}$ and $R_{\rm{ex}}$ are temperature dependent and spatially varied.
+To account for the temperature dependence of these spatially non-uniform circuit elements, we consider the device’s local effective temperature.
+(iii) The third assumption is that this local effective temperature is proportional to $T_{e}(x)$.
+These assumptions allow us to use a modified transmission line model based on the distributed-element model.
+We then explain the thermal circuit elements $R_{\rm{in}}$ and $R_{\rm{ex}}$, respectively.
 $R_{\rm{in}}$ is the thermal resistance (per unit length) for heat inflow from the heat source to the temperature measurement point (qubit).
-This thermal resistance is a series of resistances of a periodic structure, which is composed of the metallic or polysilicon gates and the insulator (we show the order estimation of $R_{\rm{in}}$, see Appendix C).
-On the basis of the above discussion, we use $R_{\rm{in}}\propto 1/T_{e}^{\beta_{\rm{in}}}$, where $\beta_{\rm{in}}$ depends on the thermal conductivity in cryogenic regime and the thickness of the gate and insulator.
-For example, $\beta_{\rm{in}}\approx 1$ when the gate is dominant and $\beta_{\rm{in}}\approx 2$ when the insulator (amorphous SiO2) is dominant (see Appendix A).
+This thermal resistance is a series of resistances of a periodic structure, which is composed of the metallic or polysilicon gates and the insulator (we show the order estimation of $R_{\rm{in}}$, see Supplementary Material II).
+On the basis of the above discussion, we can write $R_{\rm{in}}(x)=a_{\rm{in}}/(T_{e}(x)+T_{\rm{base}})^{\beta_{\rm{in}}}$, where $a_{\rm{in}}$ is a constant and $\beta_{\rm{in}}$ depends on the thermal conductivity in cryogenic regime and the thickness of the gate and insulator.
+Namely, $R_{\rm{in}}(x)$ depends only on $T_{e}(x)$ due to the local effect.
+For example, $\beta_{\rm{in}}\approx 1$ when the gate is dominant and $\beta_{\rm{in}}\approx 2$ when the insulator (amorphous SiO2) is dominant (see Supplementary Material I and Refs.
+Duthil (2014); Ekin (2006); ais ).
 $R_{\rm{ex}}$ is the thermal resistance (inverse of the thermal conductance per unit length) for heat dissipation from the heat source.
-Heat is dissipated through metallic wires and metallic structures in a dilution refrigerator and is finally collected in the mixing chamber, which acts as a cooler.
-Therefore, $R_{\rm{ex}}$ depends on the structure of the whole chip and chip’s implementation form such as the circuit board mounting the chip, bonding wires, sample packaging, and cables.
-Most of these materials are metallic.
-We then use $R_{\rm{ex}}\propto 1/T_{e}^{\beta_{\rm{ex}}}$, where $\beta_{\rm{ex}}\approx 1$ considering the metallic materials are dominant (see Appendix A).
-The heat capacitance $C_{\rm{ex}}$ is crucial for real-time thermal management because it influences the temporal changes in device temperature, which is essential for understanding the dynamic temperature effects on qubits, such as heat-induced frequency shifts Takeda et al.
-(2018); Undseth et al.
-(2023); Sato and Kawahara (2024).
-It is noted that specific heat changes drastically (approximately $\propto T^{3}$) at cryogenic temperatures, according to the Debye model, as shown in Appendix A.
-In analogy to transmission line models in electrical circuits, $T_{e}$, the temperature of the thermal rod at a point that is a distance $D$ away from the heat source having heating power $P$, can be derived as follows.
-From the telegrapher’s equations, we obtain the relations ${T_{e}(x+D)=T_{e}(x)e^{-\gamma_{\rm{th}}D}}$, ${P(x+D)=P(x)e^{-\gamma_{\rm{th}}D}}$, and ${T_{e}(x)=Z_{\rm{th}}P(x)}$, where the position on the thermal rod $x$, thermal characteristic impedance $Z_{\rm{th}}$ and thermal propagation constant $\gamma_{\rm{th}}$ are introduced.
-From these relations, we can describe
-$T_{e}(x+D)=Z_{\rm{th}}e^{-\gamma_{\rm{th}}D}P(x).$ (1)
-Since $x$ is arbitrary, we will denote $T_{e}(x+D)$ simply as $T_{e}$ and $P(x)$ as $P$ from here on.
-We then describe $Z_{\rm{th}}$ and $\gamma_{\rm{th}}$ as
-$Z_{\rm{th}}$ $=\sqrt{\frac{R_{\rm{in}}}{\left(\frac{1}{R_{\rm{ex}}}+j\omega C_{\rm{ex}}\right)}}\approx\sqrt{R_{\rm{in}}R_{\rm{ex}}}\approx a^{1+\beta}/T_{e}^{\beta},$ (2)
-$\gamma_{\rm{th}}$ $=\sqrt{R_{\rm{in}}\left(\frac{1}{R_{\rm{ex}}}+j\omega C_{\rm{ex}}\right)}\approx\sqrt{\frac{R_{\rm{in}}}{R_{\rm{ex}}}},$ (3)
-where $j$ and $\omega$ are the imaginary unit and angular frequency of thermal behaviors, respectively.
-The terms $a$ and $\beta$ are constant parameters depending on the device, where $\beta=(\beta_{\rm{in}}+\beta_{\rm{ex}})/2$.
-In the first approximation in Eqs.
-(2) and (3), $\omega$ is set to zero to perform the static analysis.
-The second approximation in Eq.
-(2) is derived by the relations $R_{\rm{in}}\propto 1/T_{e}^{\beta_{\rm{in}}}$ and $R_{\rm{ex}}\propto 1/T_{e}^{\beta_{\rm{ex}}}$.
-Substituting Eq. (2) into Eq. (1), we obtain
-$T_{e}\approx ae^{-\frac{\gamma_{\rm{th}}}{1+\beta}D}P^{\frac{1}{1+\beta}}.$ (4)
-This result indicates that the electron temperature is modeled as a function of heating power $P$, the distance $D$, and three device-dependent parameters ($a$, $\gamma_{\rm{th}}$, and $\beta$).
-Hereafter, for simplicity, we use $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\approx 1$, i.e., $\beta\approx 1$.
-This approximation is valid because, in a typical QD array structure, the thickness of the insulator is relatively lower than that of the gate, justifying $\beta_{\rm{in}}\approx 1$.
-Similarly, $\beta_{\rm{ex}}\approx 1$ is valid considering that metallic materials dominate the heat dissipation path, as mentioned above.
-Therefore, we can rewrite Eq.
-(4) as
-$T_{e}\approx ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P}.$ (5)
-We note that $a$ depends on the multiple of the thermal resistances ($R_{\rm{in}}\times R_{\rm{ex}}$) from Eq.
-(2).
-On the other hand, $\gamma_{\rm{th}}$ depends on the ratio of the thermal resistances ($R_{\rm{in}}/R_{\rm{ex}}$) from Eq.
-(3).
-Furthermore, both $a$ and $\gamma_{\rm{th}}$ are independent of $T_{e}$.
-The term $2/\gamma_{\rm{th}}$ represents the characteristic length of the heat dissipation, indicating that the heating effect decreases by $1/e$.
-Thereby, the temperature independent parameters $a$ and $\gamma_{\rm{th}}$ thermally characterize the QD array structure.
-Here, we extend our model to the case of multiple heat sources.
-Typically, the electron temperature is written as $T_{e}^{\prime}=\sqrt{T_{\rm{MXC}}^{2}+T_{e}^{2}}$, where $T_{e}^{\prime}$, $T_{\rm{MXC}}$, and $T_{e}$ are the measurable temperature, the base temperature of refrigerator, and electron temperature, respectively Noah et al.
-(2024); de Kruijf et al.
-(2024).
-Consistent with this notation, we employ the following model, i.e., the total rise in electron temperature $T_{e}^{\rm{total}}$ due to multiple heat sources is calculated using the root sum square as
-$T_{e}^{\rm{total}}=\sqrt{\sum_{i=1}^{N}(\Delta T_{e}^{i})^{2}},$ (6)
-where $N$ is the number of heat sources and $\Delta T_{e}^{i}$ is the rise in electron temperature due to the $i$-th heat source.
-From Eq.
-(6), Eq.
-(5) can be rewritten under the background temperature $T_{0}$ as
-$T_{e}^{\rm{total}}=\sqrt{T_{e}^{2}+T_{0}^{2}}=ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P+P_{0}},$ (7)
-where $T_{0}=ae^{-\frac{\gamma_{\rm{th}}}{2}D}\sqrt{P_{0}}$, and $P_{0}$ is the background heating power.
-We use this equation to fit experimental results.
-Note that in our experiment, $T_{0}$ and $P_{0}$ includes the effect of lifetime broadening and charge noise broadening de Kruijf et al.
-(2024).
+Heat is dissipated through metallic wires and metallic structures in a dilution refrigerator and is finally collected in the MXC, which acts as a cooler.
+Therefore, potentially, $R_{\rm{ex}}$ depends on the structure of the whole chip and chip’s implementation form such as the circuit board mounting the chip, bonding wires, sample packaging, and cables.
+However, we assume that the temperature dependence of this resistance is mainly determined by the local effective temperature, as mentioned in the assumption (ii).
+This is because, along the heat propagation path, regions near the gate structure have high thermal resistance, while regions farther away—such as metal wiring layers, the bottom layer, and metallic structures—have relatively low thermal resistance.
+As a result, large thermal gradients are likely to form near the gate structure, whereas the more distant regions remain close to $T_{\rm{base}}$.
+This assumption holds when a good heat dissipation structure is present around the gate region.
+We then write $R_{\rm{ex}}(x)=a_{\rm{ex}}/(T_{e}(x)+T_{\rm{base}})^{\beta_{\rm{ex}}}$, where $\beta_{\rm{ex}}=$1–2, considering the composition of the local region.
+We formulate the relations between the electron temperature $T_{e}(x)$ and the heat flow $P(x)$ as following simultaneous differential equations:
+$-\frac{dT_{e}(x)}{dx}$ $=R_{\rm{in}}(x)P(x),$ (1)
+$-\frac{dP(x)}{dx}$ $=\frac{1}{R_{\rm{ex}}(x)}T_{e}(x).$ (2)
+Here, we assume that $T_{e}(x)\gg T_{\rm{base}}$ and $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\,(\equiv\beta)$, which are referred to as assumptions (iv) and (v), respectively.
+Assumption (iv) is valid when thermal effects are significant—precisely the regime in which this analysis becomes relevant.
+The validity of assumption (v) is supported when the temperature dependence of the gate and insulator is similar, or when one of the ratios dominates, both in the local region (as considered in assumption (ii)).
+On the basis of above discussion, we can reasonably set $\beta\approx 1\text{--}2$.
+Based on these assumptions, we can rewrite the equations into an analytically solvable form as follows:
+$-\frac{dT_{e}(x)}{dx}$ $=\frac{a_{\rm{in}}}{\left[T_{e}(x)\right]^{\beta}}P(x),$ (3)
+$-\frac{dP(x)}{dx}$ $=\frac{\left[T_{e}(x)\right]^{\beta+1}}{a_{\rm{ex}}}.$ (4)
+Considering boundary conditions of $P(x=0)=P_{0}\geq 0$ and $P(x=\infty)=T_{e}(x=\infty)=0$, we can obtain one of the solutions for $x\geq 0$ and $P(x)\geq 0$ as
+$T_{e}(x)$ $=\left[a_{\rm{in}}a_{\rm{ex}}(\beta+1)\right]^{\frac{1}{2(\beta+1)}}e^{-\sqrt{\frac{a_{\rm{in}}}{a_{\rm{ex}}(\beta+1)}}x}P_{0}^{\frac{1}{\beta+1}}$
+$=ae^{-\frac{x}{L_{\rm{th}}}}P_{0}^{\frac{1}{\beta+1}},$ (5)
+$P(x)$ $=P_{0}e^{-\sqrt{\frac{a_{\rm{in}}(\beta+1)}{a_{\rm{ex}}}}x}.$ (6)
+where a coefficient $a\equiv[a_{\rm{in}}a_{\rm{ex}}(\beta+1)]^{\frac{1}{2(\beta+1)}}$ and a thermal characteristic length $L_{\rm{th}}\equiv\sqrt{\frac{a_{\rm{ex}}(\beta+1)}{a_{\rm{in}}}}$ are introduced.
+The main result in this section, Eq.
+(5), indicates that the electron temperature is modeled as a function of heating power $P_{0}$, the distance $x$, and three device-dependent parameters ($a$, $L_{\rm{th}}$, and $\beta$).
+We note that the parameter $a$ depends on the multiple of the thermal resistances coefficients ($a_{\rm{in}}\times a_{\rm{ex}}$).
+On the other hand, $L_{\rm{th}}$ depends on the ratio of the thermal resistances coefficients ($a_{\rm{ex}}/a_{\rm{in}}$).
+Furthermore, both $a$ and $L_{\rm{th}}$ are independent of $T_{e}$.
+The term $L_{\rm{th}}$ represents the characteristic length of the heat dissipation, indicating that the heating effect decreases by $1/e$.
+Thereby, the temperature independent parameters $a$ and $L_{\rm{th}}$ thermally characterize the QD array structure.
+Here, background heating power is introduced to fit the experimental results using the proposed model, Eq. (5). The total electron temperature rise, denoted by $T_{e}^{\rm{total}}$, is described as
+$T_{e}^{\rm{total}}=\left(T_{e}^{\beta+1}+T_{B}^{\beta+1}\right)^{\frac{1}{\beta+1}}=ae^{-\frac{x}{L_{\rm{th}}}}\left(P_{0}+P_{B}\right)^{\frac{1}{\beta+1}},$} (7)
+where $T_{B}=ae^{-\frac{x}{L_{\rm{th}}}}P_{B}^{\frac{1}{\beta+1}}$, and $P_{B}$ represents the background heating power.
+We use this equation to fit the experimental data.
+Note that in our experiment, $T_{B}$ and $P_{B}$ include the effects of lifetime broadening and charge noise broadening de Kruijf et al.
+(2024), whereas $P_{0}\gg P_{B}$ typically holds in most cases.
+Finally, the validity of the model is confirmed for a finite and discrete structure.
+Although assumption (i) considers a semi-infinite structure, the proposed model, Eq.
+(5), also accommodates similar circuit configurations for finite systems.
+The detailed calculation method is provided in Supplementary Material III.
+These results indicate that the device structure used in the experiment remains valid, even if it deviates from the semi-infinite configuration assumed in assumption (i).
 III Experiment
@@
 In this device, we can measure the temperature at different distances from the heating point.
-The T-shaped silicon channel (green) is fabricated using silicon on insulator (SOI), and multiple polysilicon gate electrodes are formed on the T-shaped Si channel.
+The T-shaped silicon-on-insulator (SOI) channel (green) is fabricated, and multiple polysilicon gate electrodes are formed on the T-shaped Si channel.
 The horizontal channel is covered with first gates (BG0-BG3, light blue) and second gates (PG0-PG3, blue), for a total of seven gates.
 These fine-pitch gate structures are fabricated using the self-align patterning process Lee et al.
-(2020); kun .
+(2020); Kuno et al.
+(2025).
 Each gate has terminals on both sides, allowing current to flow through them by applying DC voltages through wiring connected to instruments at room temperature.
@@
 Figure 7(b) shows the cross-section between A and A’ in Fig.
-3(a).
+7(a).
 Corresponding to the model shown in Fig.
-2(a), the unit cell length of this device is $L_{\rm{cell}}=120$ nm (we also measure the device of $L_{\rm{cell}}=160$ nm, see Appendix B).
+2(a), the unit cell length of this device is $L_{\rm{cell}}=120$ nm (we also measure the device of $L_{\rm{cell}}=160$ nm).
 The silicon chip with the above structure is glued to a cryogenic printed circuit board (QBoard, Qdevil) using silver paste, and each terminal is wired using aluminum bonding wire.
-Measurements were performed using a dilution refrigerator (Proteox, Oxford Instruments) with a base temperature of 8 mK.
+Measurements are performed using a dilution refrigerator (Proteox, Oxford Instruments) with a base temperature of 8 mK.
 III.2 Measurement
-We use CBT to estimate the local temperature at the SET.
-In CBT, there are several restrictions for QD parameters for accurately measuring temperature, e.g., the tunnel rate of barriers of SET $h\Gamma$, energy spacing of the quantum levels $\Delta E$, thermal energy $k_{B}T$, and charging energy $e^{2}/C$  Beenakker (1991).
-In this experiment, we use the classical regime, i.e., $h\Gamma\ll\Delta E\ll k_{B}T\ll e^{2}/C$.
+[…]
 By measuring the current $I_{d}$ flowing through the SET while sweeping the voltage of the SGS gate $V_{g}$ and the drain $V_{d}$ [Fig.
 4(a)], we can obtain the SGS gate voltage dependence of the differential conductance $G=dI_{d}/dV_{d}$ [Fig.
-4(b)] and estimate the electron temperature around SET $T_{e}$.
+4(b)] and estimate the electron temperature around SET as $T_{e}$.
+Note that the Coulomb diamonds shown in Fig.
+4(a) and 4(b) deviate from the ideal shape, indicating the possible formation of multiple QDs near or within the SET.
+This is likely due to structural imperfections in the device, such as disorder or surface roughness in the SET channel.
+To mitigate the influence of such imperfections, we select a Coulomb diamond that is as well-isolated as possible.
+Furthermore, the bias offset in the source-drain voltage $V_{\rm{ds}}$ observed in Fig.
+4(a) and 4(b) originates from imperfections in the measurement system—specifically, an offset in the current measurement.
+This may be attributed to the Seebeck effect in the cables or fluctuations in the ground level in the source measure unit.
+However, this offset does not significantly affect the measurement results, as it is appropriately corrected by $V_{\rm{correct}}$ during data analysis, i.e., $V_{d}=V_{\rm{sd}}-V_{\rm{correct}}$.
 The differential conductance $G$ in the region where $V_{d}$ is small is Beenakker (1991)
@@
 (8).
-This measurement method is constrained by the tunnel rate and can be used in the range $k_{B}T_{e}\gg h\Gamma$, where $h\Gamma/k_{B}\sim 1$ K in this experiment.
+This measurement method is constrained by the tunnel rate and is applicable in the regime where $k_{B}T_{e}\gg h\Gamma$.
+Given that $h\Gamma/k_{B}<0.3~\mathrm{K}$ in this experiment, the condition $k_{B}T_{e}\gg h\Gamma$ implies that the method is applicable for $T_{e}\gtrsim 1~\mathrm{K}$.
 Figure 3: (a) Schematic diagram of the device structure and setup for electron temperature measurement.
@@
 If the electric field around the SET changes during the measurement, the SET conditions are affected by the electric field and the temperature measurement cannot be performed correctly.
-Therefore, we apply a voltage of $+V/2$ and $-V/2$, where $V=\sqrt{RP}$ ($R$ is the resistance of the entire gate wiring) to each terminal of the gate to allow current to flow through the gate.
+Therefore, we apply a voltage of +V/2+V/2 and V/2-V/2, where $V=\sqrt{RP_{0}}$ ($R$ is the resistance of the entire gate wiring) to each terminal of the gate to allow current to flow through the gate.
 Due to the symmetry of the wiring structure in the dilution refrigerator, we set the voltage to be approximately zero on the channel.
-This stabilizes the SET condition because the voltage condition around the SET is stable even if the heating power $P$ is changed by varying $V$, where $P=V^{2}/R$.
-We measure the resistance of each gate wiring excluding the additional resistance on the circuit board for a low pass filter and use these values to calculate $P$, considering that the polysilicon gate has the dominant resistance among all of the gate wiring.
+This stabilizes the SET condition because the voltage condition around the SET is stable even if the heating power $P_{0}$ is changed by varying $V$, where $P_{0}=V^{2}/R$.
+We measure the resistance of each gate wiring excluding the additional resistance on the circuit board for a low pass filter and use these values to calculate $P_{0}$, considering that the polysilicon gate has the dominant resistance among all of the gate wiring.
 Figure 4: (a,b) Coulomb diamond of (a) the drain current of SET $I_{d}$ and (b) the differential conductance of SET $G$.
@@
 (8) at different temperatures.
-The estimated temperatures are (c) 2.4 K and (d) 5.5 K.
-Figure 5(b) shows the experimental results of $T_{e}$ as a function of the heating power $P$ when each gate is used as a heat source.
-As expected, $T_{e}$ increases monotonically with $P$ and $1/D$.
+The estimated temperatures are (c) 2.4 $\pm~0.1$ K and (d) 5.5 $\pm~0.1$ K, where the estimated error is the standard deviation of the $T_{e}$ estimate.
+It should be noted that the increase in the conductance peak value in (c) and (d) is most likely attributed to temperature-induced changes in the tunneling rate of the SET formed by the multiple QDs.
+The bias offset of source-drain voltage $V_{\rm{ds}}$ in (a) and (b) originates from imperfections in the measurement system, as described in the main text.
+This offset is appropriately corrected in the estimation of $T_{e}$.
+Figure 5(b) shows the experimental results of $T_{e}$ as a function of the heating power denoted by $P_{0}$ and the distance between the heater and the SET denoted by $D$ when each gate is used as a heat source.
+As expected, $T_{e}$ increases monotonically with $P_{0}$ and $1/D$.
 The experimental results can be fitted using Eq.
 (5) as
-$T_{e}=A_{i}\sqrt{P+P_{0}},$ (9)
-[…]
-We measure another device of different gate pitch as shown in Appendix B.
+$T_{e}=A_{i}(P_{0}+P_{B})^{\frac{1}{\beta+1}},$} (9)
+where $(i\in\{$BG0,BG1,BG2,BG3})\}) [see the dotted curves in Fig.
+5(b)].
+The inset of Fig.
+5(b) plots $A_{i}$ and shows the best-fit curve with $A_{i}=ae^{-\frac{D}{L_{\rm{th}}}}$ from Eq.
+(5).
+We experimentally confirm that $T_{e}$ rises is proportional to $P_{0}^{\frac{1}{\beta+1}}$ as in the model, and the heating effect of distance is reduced exponentially as expected.
+However, we note that it is difficult to determine from this experiment whether the fitted curve strictly follows exponential decay.
+To clarify this point, it would be necessary to either increase the number of measurement conditions or use a setup in which the value of $\beta$ is known—i.e., independently measured in advance by an appropriate method.
+We also measure another device of different gate pitch as shown in Fig.
+5(c).
 From these results, we conclude that the proposed model successfully reproduces the experimental results.
 This demonstrates its capability to model the thermal conduction characteristics of silicon QD array structures effectively.
-Figure 5: (a) Schematic of local heaters by flowing current in barrier gates.
-The resistance of each gate excluding the resistance of the low pass filter in Qboard at cryogenic temperatures are $R_{\rm{BG0}}=78.8~{}\rm{k}\Omega$, $R_{\rm{BG1}}=78.1~{}\rm{k}\Omega$, $R_{\rm{BG2}}=70.9~{}\rm{k}\Omega$ and $R_{\rm{BG3}}=61.3~{}\rm{k}\Omega$, respectively.
-(b) Experimental results of $T_{e}$ as a function of $P$ and $D$.
-The dotted curves are the best-fit results with Eq.
-(9).
-The inset shows the best-fit results of $A$ as a function of $D$ and the dotted curve is the best-fit result with $A=ae^{-\frac{\gamma_{\rm{th}}}{2}D}$, where $a=8.5~{}\rm{K}/\sqrt{\rm{\mu W}}$ and $2/\gamma_{\rm{th}}=280~{}\rm{nm}$.
+[…]
 IV Discussion
 To consider the limitations of our model, we re-summarize the assumptions for deriving Eq.
-(4) according to the above discussions: (i) The thermal circuit of the model is an infinite periodic structure, (ii) $R_{\rm{in}}$ and $R_{\rm{ex}}$ are temperature dependent and spatially uniform, and (iii) the net effective temperature of the QD array is proportional to $T_{e}$.
-Moreover, for deriving Eq.
-(5), we also assume that the temperature dependence of $R_{\rm{in}}$ and $R_{\rm{ex}}$ are metallic ($\beta_{\rm{in}}\approx\beta_{\rm{ex}}\approx 1$).
+(5) according to the above discussions: (i) The thermal circuit of the model is a semi-infinite periodic structure.
+(ii) Both $R_{\rm{in}}(x)$ and $R_{\rm{ex}}(x)$ are temperature-dependent and spatially non-uniform, (iii) The local effective temperature of the QD array is proportional to $T_{e}(x)$.
+(iv) $T_{e}(x)\gg T_{\rm{base}}$.
+(v) $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\,(\equiv\beta)$.
+Regarding assumption (i), numerical simulations show that even a finite and discrete circuit exhibits the same dependence as Eq.
+(5) (see Supplementary Material III).
 Despite these potential sources of modeling error, the consistency between the experimental results and the model supports the validity of our simplifications.
-The model validity is further investigated using the order estimation in Appendix C.
+Device A Device B
+$L_{\rm{cell}}$ (nm) 120 160
+[…]
+$L_{\rm{th}}$ (nm) 264 500
+$R_{\rm{in}}T_{e}$ ($\rm{K^{2}/(m\cdot W)}$) $2.7\times 10^{14}$ $1.2\times 10^{14}$
+$R_{\rm{ex}}T_{e}$ ($\rm{K^{2}\cdot m/W}$) $9.5$ $15.6$
+$R^{\rm{est}}_{\rm{in}}$ ($\rm{K^{2}/(m\cdot W)}$) $4.0\times 10^{14}$ $3.5\times 10^{14}$
+Table 1: The device-dependent parameters calculated from the experimental results assuming $\beta=1$, and order estimation results at $T_{e}=1$ K (the last row).
+Table 1 presents the device-dependent parameters estimated from the experiments, where we compare devices A and B.
+To evaluate the thermal resistances, we use the relations of $R_{\rm{in}}T_{e}^{\beta}=a^{\beta+1}/L_{\rm{th}}$ and $R_{\rm{ex}}T_{e}^{\beta}=a^{\beta+1}L_{\rm{th}}/(\beta+1)$ derived from the definitions of the $a$ and $L_{\rm{th}}$.
+Here we use $\beta=1$.
+We analyze these results to the order estimation values of heat inflow resistance $R^{\rm{est}}_{\rm{in}}$ on the basis of the device structures and the thermophysical properties (see Supplementary Material II).
+We note that the order estimation of $R_{\rm{ex}}$ is difficult due to the complicated heat dissipation path.
+We have successfully quantified the thermal characteristics of two QD array structures with different gate pitches.
+The results indicate that the two devices differ significantly in the characteristic length $L_{\rm{th}}$, while showing only minor variation in the parameter $a$.
+We believe that further experimental investigations to clarify the origin of this discrepancy will contribute to the optimization of the device structure.
 We also discuss how to design the QD array structure.
 To mitigate the thermal effects on the qubits, based on Eq.
-(5), the following conditions should be met: (I) $P$ should be reduced, (II) $D$ should be increased, and (III) the device-dependent parameters $a$ and $2/\gamma_{\rm{th}}$ should be minimized.
+(5), the following conditions should be met: (I) $P_{0}$ should be reduced, (II) $D$ should be increased, and (III) the device-dependent parameters $a$ and $L_{\rm{th}}$ should be minimized.
 Conditions (I) and (II) are straightforward, i.e., the amount of heating power needs to be reduced, for example, by lowering the electrical resistance of the gate, and the qubits need to be kept appropriately distant from the heat sources.
@@
 Reducing $a$ implies that even if the heating power is high, the temperature rise is low; in other words, the structure does not easily accumulate heat.
-To reduce $a\propto\sqrt[4]{R_{\rm{in}}R_{\rm{ex}}}$, the thermal resistance of the entire structure should be minimized.
-On the other hand, reducing $2/\gamma_{\rm{th}}$ (the characteristic length) means creating a structure that makes it difficult to transport heat to the qubits.
-To reduce $2/\gamma_{\rm{th}}\approx 2\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$, the ratio of heat dissipation to heat inflow should be increased.
+To reduce $a\propto\sqrt[2(\beta+1)]{R_{\rm{in}}R_{\rm{ex}}}$, the thermal resistance of the entire structure should be minimized.
+On the other hand, reducing $L_{\rm{th}}$ means creating a structure that makes it difficult to transport heat to the qubits —i.e., thermally separating the heat source from the qubits.
+To reduce $L_{\rm{th}}\propto\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$, the ratio of heat dissipation to heat inflow should be increased.
 For example, an additional metallic heat sink at the top and bottom of the gate layer can be effective.
@@
 (2023); Sato and Kawahara (2024).
-This will be modeled by the heat capacitance $C_{\rm{ex}}$ and measured by a high-speed readout method such as reflectometry Champain et al.
+This will be modeled by introducing the heat capacitance in the thermal circuit and measured by a high-speed readout method such as reflectometry Champain et al.
 (2024).
+The simplest case of such a dynamic thermal effect is the heating caused by gate pulsing.
+In this case, the introduction of heat capacitance leads to a thermal time constant, defined as the product of local thermal resistance and local heat capacitance on the thermal circuit, which governs the transient thermal response.
+If the time constant is shorter than the pulse duration, the thermal behavior is expected to resemble the static case, and the thermal coupling can still be characterized by $L_{\rm{th}}$.
+Conversely, if the time constant is longer than the pulse duration, the thermal impact is expected to be reduced.
+This behavior is analogous to a low-pass filter response in electrical circuits.
+Another compelling topic is phonon engineering Maire et al.
+(2017).
+If we can design a device structure in which phonons propagate coherently as waves, it may be possible to achieve more sophisticated thermal control.
+We intend to further investigate this concept and consider how our proposed model could be situated within the broader context of phonon engineering.
 In conclusion, we proposed a simple thermal circuit model for the silicon QD array structure and validated our model in experiments.
 Our proposed model is intuitive, simple, and scalable and is applicable to the wide spread of the QD array structures for thermal analysis.
+Supplementary Material
+The Supplementary Material provides additional information to support the discussion in the main text.
 Acknowledgements.
 We acknowledge N. Mertig and A. Ramsay at Hitachi Cambridge Laboratory for helpful discussions. This work is supported by JST Mooshot R&D Grant No. JPMJMS2065.
+AUTHOR DECLARATIONS
+Conflict of interest
+The authors have no conflicts to disclose.
+Data Availability
+The data that support the findings of this study are available from the corresponding author upon reasonable request.
 Appendix A Thermophysical properties
-We summarize thermophysical properties of represented materials in the typical silicon fabrication process based on Refs. Duthil (2014); Ekin (2006); ais , as shown in Fig. 6.
+We summarize thermophysical properties of represented materials in the typical silicon fabrication process based on Refs. Duthil (2014); Ekin (2006); ais , as shown in Supplementary Fig. 6.
 Figure 6: Thermophysical properties of represented materials where $\kappa$ of metallic materials (W, Cu, Al, and Poly-Si) are extrapolated by linear function ($\propto T$) in the cryo-temperature regime Duthil (2014); Ekin (2006).
 Note that Si is single crystal ais , SiO2 is amorphous Ekin (2006), and Poly-Si is P-type polysilicon (B-doped = $3\times 10^{20}$ atom/cm3ais .
-Appendix B Measurement result of device B
-We measure another device with $L_{\rm{cell}}=160$ nm (device B), similar to the device described in the main text (device A).
-Figure 7 shows the results, which exhibit the same trend as those in Fig.
-5.
-However, the electron temperature is lower than that of device A.
-This is attributed to the larger gate pitch, which implies a longer gate length and a higher ratio of the polysilicon gate to the insulator.
-This configuration likely decreases the resistance $R_{\rm{in}}$ while maintaining $R_{\rm{ex}}$, resulting in an increased thermal characteristic length $2/\gamma_{\rm{th}}\approx 2\sqrt{R_{\rm{ex}}/R_{\rm{in}}}$.
-On the other hand, $a\propto\sqrt[4]{R_{\rm{in}}R_{\rm{ex}}}$ changes minimally due to the fourth root.
-This measurement further validates the proposed thermal model.
-In Appendix C, we further summarize the experimental results.
-Figure 7: Experimental results of $T_{e}$ as a function of $P$ and $D$ in device B of $L_{\rm{cell}}=160$ nm.
-The dotted curves are the best-fit results with Eq.
-(9).
-The inset shows the best-fit results of $A$ as a function of $D$ and the dotted curve is the best-fit result with $A=ae^{-\frac{\gamma_{\rm{th}}}{2}D}$, where $a=7.0~{}\rm{K}/\sqrt{\rm{\mu W}}$ and $2/\gamma_{\rm{th}}=700~{}\rm{nm}$.
-Appendix C Summary of experimental results and order estimation of heat inflow resistance
-Device A Device B
-$L_{\rm{cell}}$ (nm) 120 160
-$a$ (K$/\sqrt{\rm{\mu W}}$) 8.5 7.0
-$2/\gamma_{\rm{th}}$ (nm) 280 700
-$R_{\rm{in}}T_{e}$ ($\rm{K^{2}/(m\cdot W)}$) $5.1\times 10^{14}$ $1.4\times 10^{14}$
-$R_{\rm{ex}}T_{e}$ ($\rm{K^{2}\cdot m/W}$) $10$ $17$
-$R^{\rm{est}}_{\rm{in}}$ ($\rm{K^{2}/(m\cdot W)}$) $4.0\times 10^{14}$ $3.5\times 10^{14}$
-Table 1: The device-dependent parameters calculated from the experimental results, and order estimation results (the last row).
-Table 1 presents the device-dependent parameters estimated from the experiments, where we compare devices A (main text) and B (Appendix B).
-To evaluate the thermal resistances, we use the relations of $R_{\rm{in}}T_{e}=a^{2}\gamma_{\rm{th}}$ and $R_{\rm{ex}}T_{e}=a^{2}/\gamma_{\rm{th}}$ derived from Eqs.
-(2) and (3), where $\beta=1$ is used.
-We analyze these results to the order estimation values of heat inflow resistance $R^{\rm{est}}_{\rm{in}}$ on the basis of the device structures and the thermophysical properties shown in Fig.
-6 (see Appendix B).
-Note that the order estimation of $R_{\rm{ex}}$ is difficult due to the complicated heat dissipation path.
-Using the model of the QD array structure shown in Fig. 2 analogous to the series resistance, we estimate $R^{\rm{est}}_{\rm{in}}$ as
+Appendix B Order estimation of heat inflow resistance
+Using the model of the QD array structure shown in Fig. 2 in the main text analogous to the series resistance, we estimate $R^{\rm{est}}_{\rm{in}}$ as
 $R^{\rm{est}}_{\rm{in}}=\frac{1}{SL_{\rm{cell}}}\left(\frac{L_{\rm{PolySi}}}{\kappa_{\rm{PolySi}}}+\frac{L_{\rm{SiO_{2}}}}{\kappa_{\rm{SiO_{2}}}}\right),$ (10)
@@
 6, we use $\kappa_{\rm{PolySi}}=0.05~\rm{W/(m\cdot K)}$ and $\kappa_{\rm{SiO_{2}}}=0.01~\rm{W/(m\cdot K)}$ assuming around 1 K.
-We obtain the estimation results as $R^{\rm{est}}_{\rm{in}}=4.0\times 10^{14}$ for device A and $R^{\rm{est}}_{\rm{in}}=3.5\times 10^{14}$ for device B at $T_{e}=1$ K, as shown in Table 1.
+We obtain the estimation results as $R^{\rm{est}}_{\rm{in}}=4.0\times 10^{14}$ for device A and $R^{\rm{est}}_{\rm{in}}=3.5\times 10^{14}$ for device B at $T_{e}=1$ K, as shown in Table 1 in the main text.
 The lower $R_{\rm{in}}$ for device B than for device A is consistent with the experimental trend.
@@
 Additionally, the order of $R^{\rm{est}}_{\rm{in}}$ values are comparable to the experimental results at $T_{e}=1$ K, validating the formulation in Eq.
-10 at least for the order estimation.
+(10) at least for the order estimation.
 We also expect that by improving this simple model or combining it with different approaches, e.g., finite element simulations, we can develop a framework for predicting thermal characteristics more reliably and accurately.
+Appendix C Finite and descrete thermal circuit model
+We present a finite and discrete thermal circuit model, analyzed using the ABCD matrix method.
+The ABCD matrix (also known as the transmission matrix) describes the relationship between input and output temperatures (analogous to voltages) and powers (analogous to currents) in a series of two-port networks shown in Supplementary Fig.
+7(a) as
+$\begin{bmatrix}T_{e}[0]\\ P[0]\end{bmatrix}$ $=\begin{bmatrix}1+\frac{R_{\rm{in}}[1]}{R_{\rm{ex}}[1]}&R_{\rm{in}}[1]\\ \frac{1}{R_{\rm{ex}}[1]}&1\end{bmatrix}\cdots\begin{bmatrix}1+\frac{R_{\rm{in}}[n]}{R_{\rm{ex}}[n]}&R_{\rm{in}}[n]\\ \frac{1}{R_{\rm{ex}}[n]}&1\end{bmatrix}\begin{bmatrix}T_{e}[n]\\ P[n]\end{bmatrix}$
+$=\begin{bmatrix}1+\frac{a_{\rm{in}}}{a_{\rm{ex}}}&\frac{a_{\rm{in}}}{T_{e}[1]^{\beta}}\\ \frac{T_{e}[1]^{\beta}}{a_{\rm{ex}}}&1\end{bmatrix}\cdots\begin{bmatrix}1+\frac{a_{\rm{in}}}{a_{\rm{ex}}}&\frac{a_{\rm{in}}}{T_{e}[n]^{\beta}}\\ \frac{T_{e}[n]^{\beta}}{a_{\rm{ex}}}&1\end{bmatrix}\begin{bmatrix}T_{e}[n]\\ P[n]\end{bmatrix}$ (11)
+where we assume $T_{e}[i]\gg T_{\rm{base}}$ and $\beta_{\rm{in}}\approx\beta_{\rm{ex}}\,(\equiv\beta)$ similar to the assumptions in the main text.
+Here, we apply the boundary conditions: $P[0]=P_{0}$ and $P[n]=0$.
+Then, we numerically solve for $T_{e}[i]$ where $i=0,\ldots,n$, given $P_{0}$.
+Supplementary Fig.
+7(b-d) illustrate the case for $n=4$ and $\beta=2$.
+The results are well fitted by Eq.
+(5) in the main text.
+In this figure, we show the results for $\beta=2$, but we have confirmed that it works well for any value of $\beta\geq 0$.
+These results show that even in a finite and discrete circuit, the parameter dependence can be modeled by Eq.
+(5) in the main text.
+Figure 7: (a) Finite and discrete thermal circuit model, and (b–d) calculation results for the case of $n=4$, where $\beta$ is set to 2, and $a_{\rm{in}}=a_{\rm{ex}}=1$.
+(b) Linear plots showing the dependence of $T_{e}$ on $P_{0}$, and (c) the corresponding log plots.
+The colored lines represent the fitting curves based on Eq.
+(5) in the main text.
+(d) Best-fit results for $A=ae^{-x/L_{\rm{th}}}$ and $\beta$, where $x=nd$, implying $A=ae^{d/L_{\rm{th}}}e^{-n}\propto e^{-n}$.
+Appendix D Validity of CBT setup
+We evaluate the validity of the CBT setup.
+In this measurement, we use the device A with $L_{\rm{cell}}=120$ nm.
+Supplementary Fig.
+8 shows the estimated electron temperature $T_{e}$ obtained by CBT under various temperatures of the mixing chamber $T_{\rm{MXC}}$ in the dilution refrigerator.
+In the high-temperature regime above 5 K, $T_{\rm{MXC}}$ and $T_{e}$ are expected to be the same.
+As expected, the results of $T_{\rm{MXC}}=T_{e}$ were obtained, confirming that the temperature estimation using this CBT measurement setup is valid.
+For the higher temperature regime of 10–-30 K shown in Fig.
+5(b) and 5(c) in the main text, the validity is supported by extrapolation from the confirmed data.
+Figure 8: Estimated electron temperature $T_{e}$ as a function of the temperature of the mixing chamber $T_{\rm{MXC}}$. Error bar shows the standard deviation error of $T_{e}$ estimate.
+We discuss the uncertainty in extracting $T_{e}$ using CBT.
+Supplementary Fig.
+9 illustrates the relationship between the estimated $T_{e}$ and the associated error for three cases, corresponding to Fig.
+5(a) and Fig.
+5(b) in the main text, and Supplementary Fig.
+8, respectively.
+The error appears to scale proportionally with the estimated $T_{e}$, suggesting that the relative error rate, $\sigma_{T_{e}}/T_{e}$, remains approximately constant throughout this measurement.
+Figure 9: Standard deviation error of $T_{e}$ estimate $\sigma_{T_{e}}$ as a function of estimated $T_{e}$ according to (a) Fig.
+5(b) and (b) Fig.
+5(c) in the main text, and (c) Supplementary Fig.
+8, respectevely.